find the gradients
3y+x=1
2y=3x+7
4y-5x=2

Answer by Alan3354(69443)

find the gradients
3y+x=1
2y=3x+7
4y-5x=2
--------------
The gradient is the slope.
3y+x=1
Solve for y.
y = (-1/3)x + 1
Slope m = the coefficient of x
m = -1/3
============
Do the others the same way.